# Benefits of Mutual Coupling in Dynamic Metasurface Antennas for Optimizing Wireless Communications – Theory and Experimental Validation

### v1

@@ III-B Non-Linearity of Radiation Pattern Parametrization
As the mutual coupling strength $\tilde{\mu}$ increases, the upper bound on $\|\hat{\mathbf{H}}_{i}^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})\|_{2}$ increases even faster than the upper bound on $\|\hat{\mathbf{j}}^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})\|_{2}$. Our analysis hence suggests that the non-linearity in the mapping from $\mathbf{c}$ to $e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ gets stronger as the mutual coupling strength is increased.
@@ III-C Summary
We showed in this section that the upper bounds on the normalized Jacobian and Hessian of the DMA radiation pattern with respect to the DMA configuration increase with the DMA’s relative mutual coupling strength. These analytical findings suggest a fundamental trade-off between strength and simplicity of the dependence of the DMA radiation pattern on the DMA configuration. Earlier works [35, 36, 53] already provided related observations regarding this trade-off in various contexts; in the next sections, we provide exhaustive systematic numerical and experimental evidence for this trade-off in the DMA context. Intuitively, the trade-off is expected based on a multi-bounce picture [37] of wave propagation in the DMA: In a DMA with stronger mutual coupling, multi-bounce wave trajectories are more prominent; the more often a wave trajectory bounces off meta-atoms, the more it is affected by their configuration and the more intertwined the effects of the meta-atoms’ configuration on the trajectory are [36].
@@ IV Full-Wave Numerical Validation
In this section, we numerically demonstrate in full-wave simulations that an increase in the DMA mutual coupling strength results in both a stronger DMA radiation pattern sensitivity and a stronger non-linearity of the DMA radiation pattern parametrization, as well as a greater ability to tailor the DMA radiation pattern to the needs of a wireless functionality. We evaluate sensitivity and non-linearity in terms of the model-based metrics from our theoretical analysis in Sec. III as well as in terms of model-free metrics that we use for our experimental validation in Sec. V, confirming that both types of metrics show the same trends as a function of mutual coupling strength. Our numerical study is not intended to reproduce every detail of our subsequent experiment.
@@ IV-A Full-Wave Simulation Details
To consider different levels of mutual coupling strength in the DMA, we alter the lossiness of the substrate in the DMA’s cavity. We use $k_{0}(1-\jmath\gamma)$ as the substrate wavenumber, where $k_{0}$ is the wavenumber in the absence of loss, and $\gamma$ controls the level of lossiness. We found that the dependence of the mutual coupling strength on $\gamma$ is smoother than on the number of vias in the fence (the latter was used in [36]).
@@ IV-B Optimization
To assess the influence of the DMA’s mutual coupling strength on the ability to tailor its radiation pattern to the requirements of a wireless functionality, we consider the prototypical problem of channel gain enhancement. The DMA is deployed as base station, and the user equipment (UE) is at a fixed location $\mathbf{r}_{\mathrm{UE}}$ in the scene. The channel gain is then directly proportional to the intensity of the component of the DMA’s radiation pattern at the UE location and parallel to the UE antenna’s orientation: $|h(\mathbf{r}_{\mathrm{UE}},k,\mathbf{c})|^{2}\propto|e^{\mathrm{rad},k}(\mathbf{r}_{\mathrm{UE}},\mathbf{c})|^{2}$, where $h(\mathbf{r}_{\mathrm{UE}},k,\mathbf{c})\in\mathbb{C}$ is the single-input single-output wireless channel from the DMA’s port to the UE antenna’s port.
@@
1 Evaluate the objective functions $\left\{\mathcal{C}\right\}_{\mathrm{init}}$ for $512$ random DMA configurations $\left\{\mathbf{c}\right\}_{\mathrm{init}}$.
@@
3 $k\leftarrow 0$
4 $i\leftarrow 1$
5 while $k<N_{\mathrm{S}}$ do
@@ 6       $i\leftarrow i+1$
7       $\mathbf{c}^{\prime}\leftarrow\mathbf{c}_{\mathrm{curr}}$ with the $\mathrm{mod}(i,N_{\rm S})$th bit flipped.
8       $\mathcal{C}_{\rm temp}\leftarrow\mathcal{C}\left(\mathbf{c}^{\prime}\right)$.
@@
10             $\mathbf{c}_{\mathrm{curr}}\leftarrow\mathbf{c}^{\prime}$
@@ 14            $k\leftarrow k+1$
Output: Optimized DMA configuration $\mathbf{c}_{\rm curr}$ and corresponding channel gain $\mathcal{C}_{\mathrm{curr}}$.
@@ IV-C2 Radiation Pattern Sensitivity
Our solver’s implementation in an automatic differentiation framework gives us direct access to the field $e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ and its partial derivative $\mathbf{j}^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ at $48\times 48$ positions $\mathbf{r}$ (see right panel in Fig. 2). In addition, since we have access to the dipole moments, we can evaluate $\hat{\mathbf{j}}^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$. We compute these three quantities for 512 random realizations of $\mathbf{c}$. Then, we evaluate two sensitivity metrics:
@@
where the average on the right hand side is over 512 random realizations of $\mathbf{c}$ and the $48\times 48$ considered positions $\mathbf{r}$ in Eq. (23a), and additionally over the entries of $\mathbf{j}^{\mathrm{rad}}(\mathbf{r},\mathbf{c})$ in Eq. (23b). $\hat{\sigma}$ is the metric used in our theoretical analysis in Sec. III and can only be evaluated given the system model; meanwhile, $\langle\sigma\rangle$ is a model-agnostic metric whose finite-difference approximation we also evaluate experimentally in Sec. V.
@@ IV-C4 Channel Gain Enhancement
where $\mathbf{c}_{\mathrm{opt}}$ is the optimized DMA configuration identified with Algorithm 1, and the averages on the right hand side are over the $48\times 48$ considered positions $\mathbf{r}$ for the numerator and additionally over 512 random configurations $\mathbf{c}_{\mathrm{rand}}$ for the denominator.
@@ IV-D Numerical Full-Wave Results
We first inspect in Fig. 3 the influence of the substrate lossiness factor $\gamma$ on the DMA’s mutual coupling strength $\mu_{\mathrm{n}}$, as well as its variant $\tilde{\mu}$. The order of magnitude of the considered values of $\gamma$ corresponds to the losses in typical printed-circuit board (PCB) substrates. As expected, we observe a monotonic decrease of the DMA’s mutual coupling strength as the substrate lossiness increases. Within the explored range of $\gamma$, $\mu_{\mathrm{n}}$ drops by a factor of roughly four and $\tilde{\mu}<1$ such that the bounds derived in Sec. III are meaningful.
Figure 3: DMA mutual coupling strength vs substrate lossiness factor $\gamma$.
@@
In the top row of Fig. 4, we observe a positive correlation between the radiation pattern sensitivity and the mutual coupling strength. This general trend is clear in terms of both the average relative magnitude of the Jacobian (left), $\langle\sigma\rangle$, and the analytically tractable normalized sensitivity metric $\langle\hat{\sigma}\rangle$ (right). Our results confirm the validity of the bound on $\langle\hat{\sigma}\rangle$, and we observe the expected scaling of the bound on $\langle\hat{\sigma}\rangle$ with $\mu_{\mathrm{n}}$. As detailed in Appendix B, the bound is not tight because we cannot control the mutual coupling between the meta-atoms; as already mentioned, future work on a beyond-diagonal DMA will overcome this limitation and attain the bound. In the second row of Fig. 4, we see that the enhanced sensitivity under stronger mutual coupling translates into concrete benefits in wireless communications. Specifically, the achieved channel gain enhancement, $\langle\eta\rangle$, displays a clear positive correlation with the sensitivity and hence with the mutual coupling strength. We observe in Fig. 4 that roughly tripling the mutual coupling strength increases the normalized sensitivity $\langle\sigma\rangle$ by 13.6 % which translates into a 36.0 % increase in channel gain enhancement. In other words, stronger mutual coupling brings about substantial improvements in our ability to tailor the DMA radiation pattern to the needs of a desired wireless functionality.
@@
The flipside of this benefit of strong mutual coupling is seen in the lower half of Fig. 4. Stronger mutual coupling increases the non-linearity of the mapping from DMA configuration to DMA radiation pattern, as clearly revealed by all three of our monitored metrics. The average relative off-diagonal magnitudes of the Hessian (left), $\langle\theta\rangle$, as well as the analytically tractable version $\langle\hat{\theta}\rangle$ clearly increase as the mutual coupling gets stronger. Again, the validity of our bound and its scaling with $\mu_{\mathrm{n}}$ are confirmed. In direct correspondance with the larger normalized Hessians for larger $\mu_{\mathrm{n}}$, we observe a clear drop in our linearity metric $\langle\zeta\rangle$ from 17.2 dB to 10.8 dB.
@@
Figure 4: Numerical full-wave results for the trade-off between strength and simplicity of the dependence of $e^{\mathrm{rad},k}(\mathbf{c})$ on $\mathbf{c}$, governed by the DMA’s mutual coupling strength $\mu_{\mathrm{n}}$. We vary $\mu_{\mathrm{n}}$ by altering the substrate lossiness $\gamma$ (see Fig. 3). The left column displays (from top to bottom) the average sensitivity $\langle\sigma\rangle$, the average channel gain enhancement $\langle\eta\rangle$, the average normalized off-diagonal Hessian entries $\langle\theta\rangle$, and the average linearity metric $\langle\zeta\rangle$, as a function of mutual coupling strength. The right column displays the differently normalized average sensitivity $\langle\hat{\sigma}\rangle$ and Hessian entries $\langle\hat{\theta}\rangle$ together with corresponding upper bounds. Insets show zoomed-in versions.
@@ IV-E Summary
To summarize this section, the trends in terms of how a DMA’s mutual coupling strength governs the trade-off between the strength and simplicity of the dependence of the DMA radiation pattern on the DMA configuration are clearly observed, irrespective of the chosen metric. The direct translation of an improved sensitivity to an improved ability to tailor the radiation pattern to the needs of a wireless functionality is also clear. Our goal in this section was not to reproduce our subsequent experiment in every detail, but to observe the described trends with both the theoretically and experimentally convenient metrics, as we did, to bridge theory and experiment.
@@ V Experimental Validation
Figure 5: Photographic images of the experimental setup and close-up views of front and back of the DMA. The DMA is fed by a coaxial cable (purple) and configured via the Arduino which receives logic commands via a USB cable (black). The other three ports seen on the DMA’s back are unused positioning fiducial that are not connected to the cavity and not used in this paper. The 96 meta-atoms are seen on the DMA’s front. A horn antenna facing the DMA is mounted on a 2D positioning stage to measure the field radiated by the DMA.
In this section, we experimentally demonstrate that an increase in the DMA mutual coupling strength results in both a stronger DMA radiation pattern sensitivity and a stronger non-linearity of the DMA radiation pattern parametrization, as well as a greater ability to tailor the DMA radiation pattern to the needs of a wireless functionality.
Our experiments are based on a single-feed chaotic-cavity-backed DMA similar to the one proposed in [16] (details below). Fabricating tens of distinct DMA prototypes that differ in the cavity’s lossiness or leakiness to systematically explore different DMA mutual coupling strengths is infeasible. Instead, we introduce an original time-gating technique that enables post-processing adjustments to the DMA’s effective mutual coupling strength. Time gating is an established post-processing technique for suppressing scattered waves in imaging applications [55, 56, 57] and antenna characterization [58, 59]. The fundamental insight is that the delay time of a path is directly related to the number of scattering events along the path – in line with the multi-bounce picture of wave propagation in the DMA. Wave energy exiting the DMA after longer delay times has undergone more bounces between the meta-atoms. If there was no mutual coupling in the DMA, there would be no wave energy exiting the DMA after long delay times. Hence, by considering only the wave energy radiated by the DMA up to some delay time $\tau$, we can suppress the contributions of paths associated with delay times beyond $\tau$, thereby effectively reducing the DMA’s effective mutual coupling strength. In other words, we have continuous control in post-processing over the DMA’s mutual coupling strength by selecting $\tau$. Of course, the mutual coupling strength in the physical prototype is an upper limit for the values that we can explore (and corresponds to $\tau\rightarrow\infty$).

### v3

@@ III-B Non-Linearity of Radiation Pattern Parametrization
As the MC strength $\tilde{\mu}$ increases, the upper bound on $\|\hat{\mathbf{H}}_{i}^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})\|_{2}$ increases even faster than the upper bound on $\|\hat{\mathbf{j}}^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})\|_{2}$. Our analysis hence suggests that the non-linearity in the mapping from $\mathbf{c}$ to $e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ gets stronger as the MC strength is increased.
@@ III-C Summary
We showed in this section that the upper bounds on the normalized Jacobian and Hessian of the DMA radiation pattern with respect to the DMA configuration increase with the DMA’s relative MC strength. These analytical findings suggest a fundamental trade-off between strength and simplicity of the dependence of the DMA radiation pattern on the DMA configuration. Earlier works [35, 36, 56, 58], as well as [59] whose preprint appeared shortly after the preprint of the present paper, already provided related observations regarding this trade-off in various contexts. In the next sections, we provide exhaustive systematic numerical and experimental evidence for this trade-off in the DMA context. Intuitively, the trade-off is expected based on a multi-bounce picture [37] of wave propagation in the DMA: In a DMA with stronger MC, multi-bounce wave trajectories are more prominent; the more often a wave trajectory bounces off meta-atoms, the more it is affected by their configuration and the more intertwined the effects of the meta-atoms’ configuration on the trajectory are [36].
@@ IV Full-Wave Numerical Validation
In this section, we numerically demonstrate in full-wave simulations that an increase in the DMA MC strength results in both a stronger DMA radiation pattern sensitivity and a stronger non-linearity of the DMA radiation pattern parametrization, as well as a greater ability to tailor the DMA radiation pattern to the needs of a wireless functionality. We evaluate sensitivity and non-linearity in terms of the model-based metrics from our theoretical analysis in Sec. III as well as in terms of model-free metrics that we use for our experimental validation in Sec. V, confirming that both types of metrics show the same trends as a function of MC strength. Our numerical study is not intended to reproduce every detail of our subsequent experiment.
@@ IV-A Full-Wave Simulation Details
To consider different levels of MC strength in the DMA, we alter the lossiness of the substrate in the DMA’s cavity. We use $k_{0}(1-\jmath\gamma)$ as the substrate wavenumber, where $k_{0}$ is the wavenumber in the absence of loss, and $\gamma$ controls the level of lossiness. We found that the dependence of the MC strength on $\gamma$ is smoother than on the number of vias in the fence (the latter was used in [36]).
@@ IV-B Optimization
To assess the influence of the DMA’s MC strength on the ability to tailor its radiation pattern to the requirements of a wireless functionality, we consider the prototypical problem of enhancing the gain of a single-user channel. This metric is a commonly considered key performance indicator because it directly correlates with end-to-end communication metrics (e.g., capacity, bit error rate, or throughput). We emphasize that our goal of maximizing the channel gain is not equivalent to other conceivable goals such as maximizing directivity or minimizing beam width that are left for future work. In addition, explorations of multi-user scenarios are deferred to future work because of their complexity and specificity to choices like the noise level. The DMA is deployed as base station, and the user equipment (UE) is at a fixed location $\mathbf{r}_{\mathrm{UE}}$ in the scene. The channel gain is then directly proportional to the intensity of the component of the DMA’s radiation pattern at the UE location and parallel to the UE antenna’s orientation: $|h(\mathbf{r}_{\mathrm{UE}},k,\mathbf{c})|^{2}\propto|e^{\mathrm{rad},k}(\mathbf{r}_{\mathrm{UE}},\mathbf{c})|^{2}$, where $h(\mathbf{r}_{\mathrm{UE}},k,\mathbf{c})\in\mathbb{C}$ is the single-input single-output wireless channel from the DMA’s port to the UE antenna’s port.
@@
1Evaluate the objective functions $\left\{\mathcal{C}\right\}_{\mathrm{init}}$ for 512512 random DMA configurations $\left\{\mathbf{c}\right\}_{\mathrm{init}}$.
@@
3 $\ell\leftarrow 0$, $i\leftarrow 0$
4 while $\ell<N_{\mathrm{M}}$ do
@@ 5  $i\leftarrow i+1$
6  $\mathbf{c}_{\mathrm{temp}}\leftarrow\mathbf{c}_{\mathrm{curr}}$ with the $(\mathrm{mod}(i-1,N_{\rm M})+1)$th bit flipped.
7  $\mathcal{C}_{\rm temp}\leftarrow\mathcal{C}\left(\mathbf{c}_{\mathrm{temp}}\right)$.
@@
9     $\mathbf{c}_{\mathrm{curr}}\leftarrow\mathbf{c}_{\mathrm{temp}}$
@@ 13    $\ell\leftarrow\ell+1$
Output: $\mathbf{c}_{\rm curr}$ and $\mathcal{C}_{\mathrm{curr}}$.
@@ IV-C2 Radiation Pattern Sensitivity
Our solver’s implementation in an automatic differentiation framework gives us direct access to the field $e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ and its partial derivative $\mathbf{j}^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ at $48\times 48$ regularly spaced grid positions $\mathbf{r}$ spanning a $0.4\times 0.4\ \mathrm{m}^{2}$ square area located 1 m in front of the DMA (see right panel in Fig. 2), i.e., in the radiating near field (DMA aperture dimension: 30 cm; DMA operation frequency: 10 GHz). In addition, since we have access to the dipole moments, we can evaluate $\hat{\mathbf{j}}^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$. We compute these three quantities for 512 random realizations of $\mathbf{c}$. Then, we evaluate two sensitivity metrics:
@@
where the average on the right hand side is over 512 random realizations of $\mathbf{c}$ and the $48\times 48$ considered positions $\mathbf{r}$ in Eq. (24a), and additionally over the entries of $\mathbf{j}^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ in Eq. (24b). $\hat{\sigma}$ is the metric used in our theoretical analysis in Sec. III and can only be evaluated given the system model; meanwhile, $\langle\sigma\rangle$ is a model-agnostic metric whose finite-difference approximation we also evaluate experimentally in Sec. V.
@@ IV-C4 Channel Gain Enhancement
where $\mathbf{c}_{\mathrm{opt}}$ is the optimized DMA configuration identified with Algorithm 1, and the averages on the right hand side are over the $48\times 48$ considered positions $\mathbf{r}$ for the numerator and additionally over 512 random configurations $\mathbf{c}_{\mathrm{rand}}$ for the denominator. Provided the transmit power $P_{\mathrm{T}}$ and receiver noise power $P_{\mathrm{N}}$ are fixed, $\langle\eta\rangle$ equals the average signal-to-noise ratio (SNR) enhancement enabled by optimizing the DMA configuration because the SNR is proportional to the channel gain: $\mathrm{SNR}=P_{\mathrm{T}}|e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})|^{2}/P_{\mathrm{N}}$.
@@ IV-D Numerical Full-Wave Results
We first inspect in Fig. 3 the influence of the substrate lossiness factor $\gamma$ on the DMA’s MC strength $\mu_{\mathrm{n}}$, as well as its variant $\tilde{\mu}$. The order of magnitude of the considered values of $\gamma$ corresponds to the losses in typical printed-circuit board (PCB) substrates. As expected, we observe a monotonic decrease of the DMA’s MC strength as the substrate lossiness increases. Within the explored range of $\gamma$, $\mu_{\mathrm{n}}$ drops by a factor of roughly four and $\tilde{\mu}<1$ such that the bounds derived in Sec. III are meaningful.
Figure 3: DMA MC strength vs substrate lossiness factor $\gamma$.
@@
In the top row of Fig. 4, we observe a positive correlation between the radiation pattern sensitivity and the MC strength. This general trend is clear in terms of both the average relative magnitude of the Jacobian (left), $\langle\sigma\rangle$, and the analytically tractable normalized sensitivity metric $\langle\hat{\sigma}\rangle$ (right). Our results confirm the validity of the bound on $\langle\hat{\sigma}\rangle$, and we observe the expected scaling of the bound on $\langle\hat{\sigma}\rangle$ with $\mu_{\mathrm{n}}$. As detailed in Appendix B, the bound is not tight because we cannot control the MC between the meta-atoms; as already mentioned, a beyond-diagonal DMA [57] can overcome this limitation and potentially attain the bound. In the second row of Fig. 4, we see that the enhanced sensitivity under stronger MC translates into concrete benefits in wireless communications. Specifically, the achieved channel gain enhancement, $\langle\eta\rangle$, displays a clear positive correlation with the sensitivity and hence with the MC strength. We observe in Fig. 4 that roughly tripling the MC strength increases the normalized sensitivity $\langle\sigma\rangle$ by 13.6 % which translates into a 36.0 % increase in channel gain enhancement. In other words, stronger MC brings about substantial improvements in our ability to tailor the DMA radiation pattern to the needs of a desired wireless functionality.
@@
The flipside of this benefit of strong MC is seen in the lower half of Fig. 4. Stronger MC increases the non-linearity of the mapping from DMA configuration to DMA radiation pattern, as clearly revealed by all three of our monitored metrics. The average relative off-diagonal magnitudes of the Hessian (left), $\langle\theta\rangle$, as well as the analytically tractable version $\langle\hat{\theta}\rangle$ clearly increase as the MC gets stronger. Again, the validity of our bound and its scaling with $\mu_{\mathrm{n}}$ are confirmed. In direct correspondence with the larger normalized Hessians for larger $\mu_{\mathrm{n}}$, we observe a clear drop in our linearity metric $\langle\zeta\rangle$ from 17.2 dB to 10.8 dB.
@@
Figure 4: Numerical full-wave results for the trade-off between strength and simplicity of the dependence of $e^{\mathrm{rad},k}(\mathbf{c})$ on $\mathbf{c}$, governed by the DMA’s MC strength $\mu_{\mathrm{n}}$. We vary $\mu_{\mathrm{n}}$ by altering the substrate lossiness $\gamma$ (see Fig. 3). The left column displays (from top to bottom) the average sensitivity $\langle\sigma\rangle$, the average channel gain enhancement $\langle\eta\rangle$, the average normalized off-diagonal Hessian entries $\langle\theta\rangle$, and the average linearity metric $\langle\zeta\rangle$, as a function of MC strength. The right column displays the differently normalized average sensitivity $\langle\hat{\sigma}\rangle$ and Hessian entries $\langle\hat{\theta}\rangle$ together with corresponding upper bounds. Insets show zoomed-in versions.
@@ IV-E Summary
To summarize this section, the trends in terms of how a DMA’s MC strength governs the trade-off between the strength and simplicity of the dependence of the DMA radiation pattern on the DMA configuration are clearly observed, irrespective of the chosen metric. The direct translation of an improved sensitivity to an improved ability to tailor the radiation pattern to the needs of a wireless functionality is also clear. Our goal in this section was not to reproduce our subsequent experiment in every detail, but to observe the described trends with both the theoretically and experimentally convenient metrics, as we did, to bridge theory and experiment.
@@ V Experimental Validation
Figure 5: Photographic images of the experimental setup and close-up views of front and back of the DMA. The DMA is fed by a coaxial cable (purple) and configured via the Arduino which receives logic commands via a USB cable (black). The other three ports seen on the DMA’s back are unused positioning fiducials that are not connected to the cavity and not used in this paper. The 96 meta-atoms are seen on the DMA’s front. A horn antenna facing the DMA is mounted on a 2D positioning stage to measure the field radiated by the DMA.
In this section, we demonstrate experimentally that an increase in the DMA MC strength results in both a stronger DMA radiation pattern sensitivity and a stronger non-linearity of the DMA radiation pattern parametrization, as well as a greater ability to tailor the DMA radiation pattern to the needs of a wireless functionality. Our objective in this section is to demonstrate a qualitative agreement between our experimental observations and our theoretical and numerical results from Sec. III and Sec. IV, respectively. Establishing a quantitative link between the experiment and our theoretical and numerical results is deferred to future work because it requires the development of techniques to estimate the model parameters for a given experimental DMA prototype, which is beyond the scope of this paper.
Our experiments are based on a single-feed chaotic-cavity-backed DMA similar to the one proposed in [15] (details below). Fabricating tens of distinct DMA prototypes that differ in the cavity’s lossiness or leakiness to systematically explore different DMA MC strengths is infeasible. Instead, we introduce an original time-gating technique that enables post-processing adjustments to the DMA’s effective MC strength. Time gating is an established post-processing technique for suppressing scattered waves in imaging applications [60, 61, 62] and antenna characterization [63, 64]. The fundamental insight is that the delay time of a path is directly related to the number of scattering events along the path – in line with the multi-bounce picture of wave propagation in the DMA. Wave energy exiting the DMA after longer delay times has undergone more bounces between the meta-atoms. If there was no MC in the DMA, there would be no wave energy exiting the DMA after long delay times. Hence, by considering only the wave energy radiated by the DMA up to some delay time $\tau$, we can suppress the contributions of paths associated with delay times beyond $\tau$, thereby effectively reducing the DMA’s effective MC strength. In other words, we have continuous control in post-processing over the DMA’s MC strength by selecting $\tau$. Of course, the MC strength in the physical prototype is an upper limit for the values that we can explore (and corresponds to $\tau\rightarrow\infty$).